# Perspective on Moreau–Yosida Regularization in Density-Functional Theory

### v1

@@ I Introduction
This perspective article aims at showing the current state of the field as well as mapping out several directions for continued research that we deem promising. We believe that, as the theory matures, it can establish itself as a mathematically rigorous but also computationally useful extension of DFT. Ideally, this means offering new and enhanced approximations for electronic-structure calculations, while also covering a wider range of physical phenomena. These and many more instances of MY regularization in DFT are discussed in this perspective article, where we also make the relevant connections to the literature.
@@
Since the original intention of introducing MY regularization into DFT was to make the theory mathematically sound, it is necessary to introduce a certain amount of mathematical concepts. We try to present this in the preliminary Section II with minimal requirements, but the interested reader might want to follow up on certain details by consulting the vast mathematical literature on the fields that we mainly rely on. These are functional and convex analysis formulated on Banach and Hilbert spaces, as well as the more specialized theory of Sobolev spaces, and the reference list includes our favorite textbooks [5; 6; 7; 8; 9; 10; 11]. After covering the most important mathematical topics in Sections II.1, II.2, and II.3, we continue in Section II.4 with a formalized presentation of DFT that introduces the universal density functional for a fully interacting system and an auxiliary, non-interacting system. This is complemented by Section II.5 on the Kohn–Sham (KS) method that we introduce with an additional offset potential to capture a priori information such as external binding potentials. Section III is concerned with introducing MY regularization into DFT in three different ways—namely, based on the universal density functional, based on the total energy functional, and based on density–potential mixing. The next two sections describe the main use cases for MY regularization in DFT to date, a technique for density–potential inversion in Section IV and a rigorous formulation of the KS method that takes a central role in the application of DFT in Section V. Section VI shows how certain formulations of DFT, such as Hartree theory and Maxwell–Schrödinger DFT, get automatically regularized through the inclusion of a mean-field energy contribution. Section VII deals with the numerical realization of density–potential inversion in the KS setting for periodic systems. Finally, we conclude with an extended outlook in Section VIII.
@@ II.1 Functional analysis preliminaries
Our development of MY regularization applied to DFT relies heavily on topological properties of the involved density and potential spaces, the Banach space $X$ and its topological dual $X^{*}$. We keep them as abstract spaces for most of our presentation, since this allows the application of the results to a huge variety of different settings. It will be indispensable to repeat important results from functional analysis, especially from the theory of Banach spaces, since this will be the first pillar on which we base DFT mathematically. The properties of the spaces $X$ and $X^{*}$ are collected in three different assumption sets of increasing restrictiveness to which later results will refer.
The basic setting for the whole presentation is a real Banach space $X$ with the norm denoted as $\|\cdot\|_{X}$. Banach spaces have been developed from the beginning as spaces of functions [12] and as such typically are infinite dimensional, which complicates their mathematical study. The feature that makes a normed vector space a Banach space is that it is complete, meaning that every Cauchy sequence in $X$ has a well-defined limit within the space $X$. Convergence to a limit value usually means strong convergence—that is, convergence in norm $\|\cdot\|_{X}$, where $x_{n}\to x$ in $X$ if and only if
@@
$X^{*}=\{u:X\to\mathbb{R}\mid u\;\text{linear and continuous}\}.$ (2)
Now, $X^{*}$ gets equipped with a norm itself, the operator norm on $X$, that we will accordingly denote with $\|\cdot\|_{X^{*}}$,
@@
The completeness of $X$ then carries over to $X^{*}$, making it a Banach space as well. With the help of $X^{*}$ we can define weak convergence $x_{n}\rightharpoonup x$ in $X$ as the property that, for all $u\in X^{*}$, it holds that
@@
be a set-valued map that maps each element in $X$ to a subset of $X^{*}$, called the duality map. The Hahn–Banach theorem guarantees that the image of every $x\in X$ is non-empty and it is easy to see that $J$ is homogeneous (i.e., $J(\alpha x)=\alpha J(x)$ for every $\alpha\in\mathbb{R}$), but we would like to impose further topological properties on $X$ so as to make the duality mapping nicer. A Banach space for which $J(x)$ is always single-valued (and thus has not to be considered as a set-valued map) is called smooth. In a smooth Banach space, the duality map is unique and given by the Gâteaux (functional) derivative of the convex function $\phi(x)=\frac{1}{2}\|x\|_{X}^{2}$,
@@
Interestingly, the smoothness of a space $X$ implies that its dual $X^{*}$ is strictly convex (the line segment between any two points $u\neq v$, $\|u\|_{X^{*}}=\|v\|_{X^{*}}=1$, lies strictly within the unit ball). This is an expression of the general connection of smoothness properties of a Banach space to convexity features on the dual side. The next important property for our space $X$ is that of reflexivity, which means that the bidual $X^{**}=(X^{*})^{*}$ is isometrically isomorphic to the space $X$ itself and can be identified with it. Hence, if $X$ is reflexive, then $X$ can be recovered from $X^{\ast}$ as $X=X^{**}=(X^{*})^{*}$.
If $X$ is reflexive, then $X^{*}$ is also reflexive. We may therefore set up a duality map $J^{\ast}$ from $X^{\ast}$ to $X=X^{\ast\ast}$. For every $u\in X^{*}$, we can consider an $x\in J^{*}(u)=\{x\in X\mid u(x)=\|x\|_{X}^{2}=\|u\|_{X^{*}}^{2}\}$ and with this recover $u\in J(x)$, so $J$ is also surjective if $X$ is reflexive. For a surjective function we can consider its set-valued inverse and we thus get $J^{-1}=J^{*}$. Now note that for $X^{*}$ also smooth ($X$ strictly convex) we have $J^{-1}$ single-valued as well, so $J$ is indeed bijective. We collect all this properties for the minimal assumptions for our Banach spaces $X$ and $X^{*}$:
@@
The real Banach space $X$ is reflexive, with $X$ and $X^{*}$ strictly convex. The corresponding duality map $J:X\to X^{*}$ is single-valued with a well-defined inverse $J^{-1}:X^{*}\to X$.
The prototypical example for such Banach spaces, which is also relevant in the context of DFT, are the Lebesgue spaces $L^{p}$, $1<p<\infty$. It is important to note that $p=1$ and $p=\infty$ must be excluded to fulfill Assumption 1, since those spaces are neither reflexive nor strictly convex. At this point, it already becomes obvious that we need to adapt the typical choice of density space in DFT [3], $L^{1}(\mathbb{R}^{3})\cap L^{3}(\mathbb{R}^{3})$, since by including $L^{1}$ the space becomes non-reflexive and thus cannot fulfill Assumption 1. It is precisely for this reason that, for example, Eschrig [13] in his discussion of thermal DFT switches to the reflexive space $L^{3}(\Omega)$ on a bounded three-dimensional periodic domain, which anyway entails $L^{3}(\Omega)\subset L^{1}(\Omega)$.
Assumption 1 has been our standard assumption in several previous publications on Moreau–Yosida regularization in DFT [14; 15; 4] and is sufficient for many use cases. Yet, in some instances, additional regularity is required and we add the condition that $X^{*}$ is also uniformly convex. Note that uniform convexity implies reflexivity. A space $X^{*}$ is called uniformly convex if, for each $\varepsilon\in(0,2]$, there is a $\delta(\varepsilon)>0$ such that, for all $u,v\in X^{*}$ with $\|u\|_{X^{*}}=\|v\|_{X^{*}}=1$ and $\|u-v\|_{X^{*}}=\varepsilon$, we have $1-\frac{1}{2}\|u+v\|_{X^{*}}\geq\delta(\varepsilon)$. Adding the requirement of uniform convexity, we obtain:
@@
$X$ and $X^{*}$ fulfill all requirements of Assumption 1 and additionally $X^{*}$ is assumed uniformly convex. The duality map $J:X\to X^{*}$ is then single-valued and uniformly continuous.
Finally, we can tighten those assumptions even more and move to a Hilbert space $X$, equipped with an inner product denoted as $\langle\cdot,\cdot\rangle_{X}$. Hilbert spaces are automatically uniformly convex and thus also reflexive. Since then also $X^{*}$ is a Hilbert space, we have all the properties in Assumption 2 fulfilled. For Hilbert spaces, the duality map $J:X\to X^{*}$ is exactly the isometric isomorphism from the Riesz representation theorem and linear as such. Nevertheless, we do not want to identify a Hilbert space with its dual in the given context. The original work introducing Moreau–Yosida regularization to DFT of Kvaal et al. [1] considered a Hilbert-space setting and it is also employed elsewhere [16; 17; 18; 19]. We also highlight a comprehensive chapter by Kvaal [20] in a recent textbook on DFT.
@@
The Banach space $X$ and consequently also $X^{*}$ are Hilbert spaces. The duality map $J:X\to X^{*}$ is then linear and Lipschitz continuous with constant 1. Even though $X=X^{*}$, we will not assume $J$ chosen to identify the space with its dual.
@@
This construction can be generalized to arbitrary $X$ and $X^{*}$ within the formalism of a rigged Hilbert space (Gel’fand triple). Applied to the Hilbert space of wavefunctions, this approach provides a rigorous foundation for Dirac’s bra–ket formalism and can be employed to describe continuous spectra and scattering states in quantum mechanics [21; 22]. Those are different objects than the densities and potentials considered here, but we find the similarities still striking.
@@ II.2 Convex analysis preliminaries
The second mathematical pillar of DFT is without any doubt convex analysis, since DFT admits a beautiful formulation in terms of convex and concave functionals on Banach spaces. On a Banach space $X$ that fulfills Assumption 1, we now consider convex functions $f\colon X\to\mathbb{R}\cup\{+\infty\}$ that are not $\infty$ everywhere. We also usually assume lower semicontinuity, which means that for each $x_{0}\in X$ it holds that $\liminf_{x\to x_{0}}f(x)\geq f(x_{0})$. Since the argument $x\in X$ of $f(x)$ is the element of a Banach space of functions that gets mapped to the scalar field, $f$ is sometimes called a functional (especially in the context of DFT). The Legendre–Fenchel transform (convex conjugate) of any $f:X\to\mathbb{R}\cup\{-\infty,+\infty\}$ is a function $f^{*}\colon X^{*}\to\mathbb{R}\cup\{-\infty,+\infty\}$ given by
@@
The convex conjugate of a convex and lower semicontinuous $f\colon X\to\mathbb{R}\cup\{+\infty\}$ is again convex and lower semicontinuous. By applying this transformation twice and restricting the result again to $X\subseteq X^{**}$ (biconjugate)
$f^{**}(x)=\sup_{u\in X^{*}}\{\langle u,x\rangle-f^{*}(u)\}$ (10)
@@
On the real axis, the set $\partial f(x)$ can be interpreted as the collection of all slopes for which the corresponding tangent of $f$ at $x$ never exceeds the function $f$ itself (globally, since the function is assumed convex). It can be defined for concave functions as the superdifferential by just reverting the inequality, and we use the same notation in this case. Note that $\partial f(x)$ is always a convex set, and that it includes infinitely many elements where $f$ has a kink and is thus non-differentiable. It can also be the empty set. To illustrate this, take $X=\mathbb{R}$ and $f(x)=-\sqrt{x}$ for $x\geq 0$ and $f(x)=+\infty$ for $x<0$ (which actually connects to reduced-density-matrix functional theory [23]) and then note that $\partial f(0)=\emptyset$. Another one-dimensional example can be taken from a DFT-treatment of the quantum Rabi model [24], where $\partial f(\pm 1)=\emptyset$ for $f(x)=-\sqrt{1-x^{2}}$ for $x\in[-1,1]$ and $f(x)=+\infty$ outside. If, on the other hand, $f$ is differentiable at $x$, then the subdifferential $\partial f(x)$ is single-valued and consists only of the functional derivative $\nabla f(x)$. The subdifferential fulfills the usual role of a derivative, by qualifying the minimum of a convex (not necessarily lower semicontinuous) functional,

### v2

@@ I Introduction
This perspective article aims at showing the current state of the field as well as mapping out several directions for continued research that we deem promising. We believe that, as the theory matures, it can establish itself as a mathematically rigorous but also computationally useful extension of DFT. Ideally, this means offering new and enhanced approximations for electronic-structure calculations, while also covering a wider range of physical phenomena. A main result of this work (discussed in Section II.4) is that the necessary function-space framework for MY regularization can be combined with the usual potential space of quantum mechanics only if the setting is such that it always guarantees existence of a ground state. Another interesting finding, laid out in Section III.3, is the possible connection of the topology to physics, in that the norm of the density and potential spaces exactly gives their energy content and that they are connected via the Poisson equation. These and many more insights into MY regularization applied to DFT are discussed in this perspective article, where we also make the relevant connections to the literature.
@@
Since the original intention of introducing MY regularization into DFT was to make the theory mathematically sound, it is necessary to introduce a certain amount of mathematical concepts. We try to present this in the preliminary Section II with minimal requirements, but the interested reader might want to follow up on certain details by consulting the vast mathematical literature on the fields that we mainly rely on. These are functional and convex analysis formulated on Banach and Hilbert spaces, as well as the more specialized theory of Sobolev spaces, and the reference list includes our favorite textbooks [8; 9; 111; 85; 26; 10; 1; 104]. After covering the most important mathematical topics in Sections II.1, II.2, and II.3, we continue in Section II.4 with a formalized presentation of DFT that introduces the universal density functional for a fully interacting system and for an auxiliary, non-interacting system. This is complemented by Section II.5 on the Kohn–Sham (KS) method, which we introduce with an additional offset potential to capture a priori information such as external binding potentials. Section III is concerned with introducing MY regularization into DFT in three different ways—namely, based on the universal density functional, based on the total energy functional, and based on density–potential mixing. The next two sections describe the main use cases for MY regularization in DFT to date: a technique for density–potential inversion in Section IV and a rigorous formulation of the KS method that takes a central role in the application of DFT in Section V. Section VI shows how certain formulations of DFT, such as Hartree theory and Maxwell–Schrödinger DFT, get automatically regularized through the inclusion of a mean-field energy contribution. Section VII deals with the numerical realization of density–potential inversion in the KS setting for periodic systems. Finally, we conclude with an extended outlook in Section VIII. We provide a brief notational aid for the most frequently used symbols in Table 1.
Table 1: Notational aid for frequently used symbols.
[…]
$X^{*}$ generic Banach space for potentials Sec. II.4
$D$ set of $N$-representable densities Eq. (II.4)
$J$ duality map from $X$ to $X^{*}$ Eq. (5)
$f^{*}(u)$ Legendre–Fenchel transform Eq. (9)
(convex conjugate)
$\partial f(x)$ sub/superdifferential of a Eq. (11)
convex/concave functional
$f_{\varepsilon}(x)$ MY regularization Eq. (15)
$\nabla f_{\varepsilon}(x)$ functional derivative Sec. II.3
$\mathop{\mathrm{prox}}_{\varepsilon f}(x)$ proximal map Eq. (16)
$F^{\lambda}(\rho)$ universal functional Eq. (31)
$E^{\lambda}(v)$ ground-state energy functional Eq. (24)
$F_{\varepsilon}^{\lambda}(x)$ MY-regularized universal functional Eq. (49)
$E_{\varepsilon}^{\lambda}(v)$ convex conjugate of $F_{\varepsilon}^{\lambda}(x)$ Eq. (50)
@@ II Preliminaries
The following five subsections introduce the mathematical and theoretical prerequisites for performing MY regularization in DFT and for studying related topics. Section II.1 deals with some basic notions of functional analysis, most importantly certain properties of Banach spaces that are later required for MY regularization as well as defining the duality map. Section II.2 gives some central results needed for the convex analysis of DFT. Section II.3 is the last mathematical part and introduces MY regularization with its most important properties. Section II.4 continues with a presentation of formal DFT and contains an important result that connects the properties of the function spaces with the existence of ground states. Finally, Section II.5 briefly introduces the KS method of DFT.
@@ II.1 Functional analysis preliminaries
Our development of MY regularization applied to DFT relies heavily on topological properties of the involved density and potential spaces, the Banach space $X$ and its topological dual $X^{*}$. We keep them as abstract spaces for most of our presentation, since this allows the application of the results to a huge variety of different settings. It will be indispensable to repeat important results from functional analysis, especially from the theory of Banach spaces, since this will be the first pillar on which we base DFT mathematically. The properties of the spaces $X$ and $X^{*}$ are collected in three different assumption sets of increasing restrictiveness, to which later results will refer.
The basic setting for the whole presentation is a real Banach space $X$ with the norm denoted as $\|\cdot\|_{X}$. Banach spaces have been developed from the beginning as spaces of functions [7] and as such typically are infinite dimensional, which complicates their mathematical study. The feature that makes a normed vector space a Banach space is that it is complete, meaning that every Cauchy sequence in $X$ has a well-defined limit within the space $X$. Convergence to a limit value usually means strong convergence—that is, convergence in the norm $\|\cdot\|_{X}$, where $x_{n}\to x$ in $X$ if and only if
@@
$X^{*}=\{u\colon X\to\mathbb{R}\mid u\;\text{linear and continuous}\}.$ (2)
Now, $X^{*}$ gets equipped with a norm itself, the operator norm on $X$, which we will accordingly denote with $\|\cdot\|_{X^{*}}$,
@@
The completeness of $X$ then carries over to $X^{*}$, making it a Banach space as well. With the help of $X^{*}$, we can define weak convergence $x_{n}\rightharpoonup x$ in $X$ as the property that, for all $u\in X^{*}$, it holds that
@@
be a set-valued map that maps each element in $X$ to a subset of $X^{*}$, called the duality map. This generalizes the Riesz representation for Hilbert spaces, that is a one-to-one map, to the case of Banach spaces. The Hahn–Banach theorem guarantees that the image of every $x\in X$ under $J$ is non-empty and it is easy to see that $J$ is homogeneous (i.e., $J(\alpha x)=\alpha J(x)$ for every $\alpha\in\mathbb{R}$), but we would like to impose further topological properties on $X$ so as to make the duality mapping nicer. A Banach space for which $J(x)$ is always single-valued (and thus has not to be considered as a set-valued map) is called smooth. In a smooth Banach space, the duality map is unique and given by the Gâteaux (functional) derivative of the convex function $\phi(x)=\frac{1}{2}\|x\|_{X}^{2}$,
@@
Interestingly, the smoothness of a space $X$ implies that its dual $X^{*}$ is strictly convex (meaning that the line segment between any two points $u\neq v$, $\|u\|_{X^{*}}=\|v\|_{X^{*}}=1$, lies strictly within the unit ball). This is an expression of the general connection of smoothness properties of a Banach space to convexity features on the dual side.
The next important property for our space $X$ is that of reflexivity, which means that the bidual $X^{**}=(X^{*})^{*}$ is isometrically isomorphic to the space $X$ itself and can be identified with it. Hence, if $X$ is reflexive, then $X$ can be recovered from $X^{\ast}$ as $X=X^{**}=(X^{*})^{*}$.
If $X$ is reflexive, then $X^{*}$ is also reflexive. We may therefore also set up a duality map $J^{\ast}$ from $X^{\ast}$ to $X=X^{\ast\ast}$. For every $u\in X^{*}$, we can consider an $x\in J^{*}(u)=\{x\in X\mid u(x)=\|x\|_{X}^{2}=\|u\|_{X^{*}}^{2}\}$ and with this recover $u\in J(x)$, so $J$ is also surjective if $X$ is reflexive. For a surjective function, we can consider its set-valued inverse and we thus get $J^{-1}=J^{*}$. Now note that, when $X^{*}$ is also smooth ($X$ strictly convex), we have $J^{-1}$ single-valued as well, so $J$ is indeed bijective [8, Prop. 1.117]. We collect all these properties as the minimal assumptions for our Banach spaces $X$ and $X^{*}$:
@@
The real Banach space $X$ is reflexive, with $X$ and $X^{*}$ strictly convex. The corresponding duality map $J\colon X\to X^{*}$ is then single-valued with a well-defined inverse $J^{-1}\colon X^{*}\to X$.
The prototypical example for such Banach spaces, which is also relevant in the context of DFT, are the Lebesgue spaces $L^{p}$, $1<p<\infty$. It is important to note that $p=1$ and $p=\infty$ are excluded by Assumption 1, since those spaces are neither reflexive nor strictly convex. At this point, it already becomes obvious that we need to adapt the typical choice of density space in DFT [63], $L^{1}(\mathbb{R}^{3})\cap L^{3}(\mathbb{R}^{3})$, since by including $L^{1}$ the space becomes non-reflexive and thus cannot fulfill Assumption 1. It is precisely for this reason that, for example, Eschrig [31] in his discussion of thermal DFT switches to the reflexive space $L^{3}(\Omega)$ on a bounded three-dimensional periodic domain, which anyway entails $L^{3}(\Omega)\subset L^{1}(\Omega)$.
Assumption 1 has been our standard assumption in several previous publications on Moreau–Yosida regularization in DFT [58; 56; 73] and is sufficient for many use cases. Yet, in some instances, additional regularity is required and we add the condition that $X^{*}$ is also uniformly convex. Note that uniform convexity implies reflexivity. A space $X^{*}$ is called uniformly convex if, for each $\varepsilon\in(0,2]$, there is a $\delta(\varepsilon)>0$ such that, for all $u,v\in X^{*}$ with $\|u\|_{X^{*}}=\|v\|_{X^{*}}=1$ and $\|u-v\|_{X^{*}}=\varepsilon$, we have $1-\frac{1}{2}\|u+v\|_{X^{*}}\geq\delta(\varepsilon)$. Adding the requirement of uniform convexity, we obtain:
@@
$X$ and $X^{*}$ fulfill all requirements of Assumption 1 and additionally $X^{*}$ is assumed uniformly convex. The duality map $J\colon X\to X^{*}$ is then additionally uniformly continuous.
Finally, we can tighten those assumptions even more and move to a Hilbert space $X$, equipped with an inner product denoted as $\langle\cdot,\cdot\rangle_{X}$. Hilbert spaces are automatically uniformly convex and thus also reflexive. Since then also $X^{*}$ is a Hilbert space, we have all the properties in Assumption 2 fulfilled. For Hilbert spaces, the duality map $J\colon X\to X^{*}$ is exactly the isometric isomorphism from the Riesz representation theorem and linear as such. Nevertheless, we do not want to identify a Hilbert space with its dual in the given context. The original work introducing Moreau–Yosida regularization to DFT of Kvaal et al. [54] considered a Hilbert-space setting and it is also employed elsewhere [18; 94; 24; 43; 11]. We also highlight a comprehensive chapter by Kvaal [55] in a recent textbook on DFT.
@@
The Banach space $X$ and consequently also $X^{*}$ are Hilbert spaces. The duality map $J\colon X\to X^{*}$ is then linear and Lipschitz continuous with constant 1, with a well-defined inverse $J^{-1}\colon X^{*}\to X$ that has the same properties.
@@
This construction can be generalized to arbitrary $X$ and $X^{*}$ within the formalism of a rigged Hilbert space (Gel’fand triple). Applied to the Hilbert space of wavefunctions, this approach provides a rigorous foundation for Dirac’s bra–ket formalism and can be employed to describe continuous spectra and scattering states in quantum mechanics [12; 66]. Those are different objects than the densities and potentials considered here, but we find the similarities still striking.
@@ II.2 Convex analysis preliminaries
The second mathematical pillar of DFT is without any doubt convex analysis, since DFT admits a beautiful formulation in terms of convex and concave functionals on Banach spaces. On a Banach space $X$ that fulfills Assumption 1, we now consider convex functions $f\colon X\to\mathbb{R}\cup\{+\infty\}$ that are not $\infty$ everywhere. We also usually assume lower semicontinuity, which means that, for each $x_{0}\in X$, it holds that $\liminf_{x\to x_{0}}f(x)\geq f(x_{0})$. Since the argument $x\in X$ of $f(x)$ is the element of a Banach space of functions that gets mapped to the scalar field, $f$ is sometimes called a functional (especially in the context of DFT). The Legendre–Fenchel transform (convex conjugate) of any $f\colon X\to\mathbb{R}\cup\{-\infty,+\infty\}$ is a function $f^{*}\colon X^{*}\to\mathbb{R}\cup\{-\infty,+\infty\}$ given by
@@
The convex conjugate of a convex and lower semicontinuous $f\colon X\to\mathbb{R}\cup\{+\infty\}$ is again convex and lower semicontinuous. By applying this transformation twice and restricting the result again to $X\subseteq X^{**}$ (biconjugate),
$f^{**}(x)=\sup_{u\in X^{*}}\{\langle u,x\rangle-f^{*}(u)\},$ (10)
@@
On the real line, the set $\partial f(x)$ can be interpreted as the collection of all slopes for which the corresponding tangent of $f$ at $x$ never exceeds the function $f$ itself (globally, since the function is assumed convex). It can be defined for concave functions as the superdifferential by just reverting the inequality, and we use the same notation in this case. Note that $\partial f(x)$ is always a convex set, and that it includes infinitely many elements where $f$ has a kink and is thus non-differentiable. It can also be the empty set. To illustrate this, take $X=\mathbb{R}$ and $f(x)=-\sqrt{x}$ for $x\geq 0$ and $f(x)=+\infty$ for $x<0$ (which actually connects to reduced-density-matrix functional theory [89]) and then note that $\partial f(0)=\emptyset$. Another one-dimensional example can be taken from a DFT-treatment of the quantum Rabi model [6], where $\partial f(\pm 1)=\emptyset$ for $f(x)=-\sqrt{1-x^{2}}$ for $x\in[-1,1]$ and $f(x)=+\infty$ outside. If, on the other hand, $f$ is differentiable at $x$, then the subdifferential $\partial f(x)$ is single-valued and consists only of the functional derivative $\nabla f(x)$. An graphical illustration with a function on $\mathbb{R}$ is provided in Figure 1.
The subdifferential fulfills the usual role of a derivative, by qualifying the minimum of a convex (not necessarily lower semicontinuous) functional,